# ExpertSteer: Intervening in LLMs through Expert Knowledge

### v1

@@ 4.2 Overall Performance
As shown in Table 2, ExpertSteer consistently boosts performance across three target models and three domains, outperforming other intervention methods and matching or surpassing fully fine-tuned approaches like SFT and KD. In the Medical and Financial domains, it provides average gains of +4.98 for Llama-3.1-8B-Instruct and +5.34 for Qwen2.5-7B-Instruct. Furthermore, ExpertSteer consistently outperforms SFT and KD in the Mathematical domain, demonstrating its superior efficiency for highly complex tasks. Even when target models occasionally outperforms expert models, ExpertSteer discovers additional knowledge through steering vectors. For example, on the FPB benchmark, the Qwen2.5-7B-Instruct baseline and expert models achieve scores of 76.23 and 64.34, respectively, while ExpertSteer achieves 78.40. This underscores the effectiveness of ExpertSteer in transferring expertise.
@@
In practice, expert and target models are likely to come from different families. Hence, we evaluate ExpertSteer under both same-family (   $\mathcal{SF}$) and cross-family (   $\mathcal{XF}$) settings, where same-family indicates that the expert model and the target model belong to the same model family, while cross-family indicates that they belong to different families. As shown in Table 2, ExpertSteer consistently outperforms the baseline in both settings, showing gains of +4.98, +5.51, and +1.34 in three domains for same-family, and +4.38 (Medical) and +5.34 (Financial) in cross-family settings using Qwen2.5-7B-Instruct as the target. These results confirm that ExpertSteer effectively extracts and transfers expertise despite model disparities, demonstrating its applicability and generalizability.
@@
Qwen2.5-7B-Instruct   $\mathcal{SF}$
@@
Gemma-2-2b-Instruct   $\mathcal{XF}$
@@
While ExpertSteer effectively transfers knowledge across domains, we also investigate its potential to enhance model performance when both the expert and target models belong to the same domain. To this end, we leverage the general-purpose Qwen2.5-14B-Instruct as the expert model and present the results in Table 3. The results demonstrate that ExpertSteer consistently outperforms other steering methods on both natural language understanding (NLU) and safety tasks. Unlike prior steering methods, which are often constrained by the model’s inherent capabilities, ExpertSteer effectively leverages the strengths of more powerful models, thereby unlocking their full potential. These findings highlight the versatility and effectiveness of ExpertSteer in both cross-domain and same-domain scenarios.
@@ 5.2 Discussion
Building on our earlier findings that ExpertSteer boosts performance, we now explore its impact on base models by applying it to MMLU tasks with Llama-3.1-8B and Qwen2.5-7B. As shown in Table 6, although ExpertSteer again improves results, the gains are smaller than with SFT target models, referring to Table 2, because the steering vectors (derived from SFT expert models) face a larger distributional gap when applied to base models. This gap reduces effectiveness of the steering vectors in transferring expertise to the base models.
@@ Appendix F Supplementary Results of Vector Generation Source
The steering vectors used in previous studies are extracted from the model itself [22, 23], but we argue that the steering vectors should be more effective if they are generated by expert models. In this section, we investigate the effectiveness of using steering vectors generated from the model itself (Self-generated) and those generated from expert models (Expert-generated). As shown in Table 9, we find that the steering vectors generated from expert models are more effective than those generated from the model itself. This indicates that the steering vectors generated from expert models can better capture additional knowledge and improve the performance of ExpertSteer. These findings are

### v2

@@ 4.2 Overall Performance
As shown in Section 4, ExpertSteer consistently boosts performance across three target models and three domains, outperforming other intervention methods and matching or surpassing fully fine-tuned approaches like SFT and KD. In the Medical and Financial domains, it provides average gains of +4.98 for Llama-3.1-8B-Instruct and +5.34 for Qwen2.5-7B-Instruct. Furthermore, ExpertSteer consistently outperforms SFT and KD in the Mathematical domain, demonstrating its superior efficiency for highly complex tasks. Even when target models occasionally outperforms expert models, ExpertSteer discovers additional knowledge through steering vectors. For example, on the FPB benchmark, the Qwen2.5-7B-Instruct baseline and expert models achieve scores of 76.23 and 64.34, respectively, while ExpertSteer achieves 78.40. This underscores the effectiveness of ExpertSteer in transferring expertise.
@@
In practice, expert and target models are likely to come from different families. Hence, we evaluate ExpertSteer under both same-family ( $\mathcal{SF}$) and cross-family ( $\mathcal{XF}$) settings, where same-family indicates that the expert model and the target model belong to the same model family, while cross-family indicates that they belong to different families. As shown in Section 4, ExpertSteer consistently outperforms the baseline in both settings, showing gains of +4.98, +5.51, and +1.34 in three domains for same-family, and +4.38 (Medical) and +5.34 (Financial) in cross-family settings using Qwen2.5-7B-Instruct as the target. These results confirm that ExpertSteer effectively extracts and transfers expertise despite model disparities, demonstrating its applicability and generalizability.
@@
gray!20        Qwen2.5-7B-Instruct      $\mathcal{SF}$
@@
gray!20        Gemma-2-2b-Instruct      $\mathcal{XF}$
@@
While ExpertSteer effectively transfers knowledge across domains, we also investigate its potential to enhance model performance when both the expert and target models belong to the same domain. To this end, we leverage the general-purpose Qwen2.5-14B-Instruct as the expert model and present the results in Section 4.2. The results demonstrate that ExpertSteer consistently outperforms other steering methods on both natural language understanding (NLU) and safety tasks. Unlike prior steering methods, which are often constrained by the model’s inherent capabilities, ExpertSteer effectively leverages the strengths of more powerful models, thereby unlocking their full potential. These findings highlight the versatility and effectiveness of ExpertSteer in both cross-domain and same-domain scenarios.
@@ 5.2 Discussion
Building on our earlier findings that ExpertSteer boosts performance, we now explore its impact on base models by applying it to MMLU tasks with Llama-3.1-8B and Qwen2.5-7B. As shown in Table 6, although ExpertSteer again improves results, the gains are smaller than with SFT target models, referring to Section 4, because the steering vectors (derived from SFT expert models) face a larger distributional gap when applied to base models. This gap reduces effectiveness of the steering vectors in transferring expertise to the base models.
@@ 6 Conclusion
Acknowledgement
This work is funded by EU Horizon Europe (HE) Research and Innovation programme grant No 101070631, and UK Research and Innovation under the UK HE funding grant No 10039436.
The computations described in this research were performed using the Baskerville Tier 2 HPC service (https://www.baskerville.ac.uk/). Baskerville was funded by the EPSRC and UKRI through the World Class Labs scheme (EP/T022221/1) and the Digital Research Infrastructure programme (EP/W032244/1) and is operated by Advanced Research Computing at the University of Birmingham.
@@ Appendix F Supplementary Results of Vector Generation Source
black
The steering vectors used in previous studies are extracted from the model itself [22, 23], but we argue that the steering vectors should be more effective if they are generated by expert models. In this section, we investigate the effectiveness of using steering vectors generated from the model itself (Self-generated) and those generated from expert models (Expert-generated). As shown in Appendix F, we find that the steering vectors generated from expert models are more effective than those generated from the model itself. This indicates that the steering vectors generated from expert models can better capture additional knowledge and improve the performance of ExpertSteer. These findings are